OpenStudy (anonymous):

If x is a whole number and 437 = (21 + x)(21 - x), then x =

OpenStudy (anonymous):

437 = (21+x)(21-x) Foil (21+x)(21-x)

OpenStudy (anonymous):

441?

OpenStudy (anonymous):

(21+x)(21-x) = (21*21)+(21*-x)+(x*21)+(x*-x)

OpenStudy (anonymous):

(21*x)+(21*-x) cancel since 21x+(-21x) = 0

OpenStudy (paki):

437 = (21 + x)(21 - x) 441-21x+21x-x^2= 437 x^2=441-437 x^2=4 x=2...

OpenStudy (anonymous):

oh ok, Thank you

OpenStudy (paki):

pleasure